Grade upto college level Physical Chemistry

The bond dissociation energies of gaseous H2, CI2 and HCI are 104, 58 and 103 kcal/mole respectively. Calculate the enthyalpy of formation of HCI gas.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Bond H – H CI – CI H – CI
∆H disso. 104 kcal 58 kcal 103 kcal
Formation of hydrogen chloride can be represented as
H – H + CI – CI → 2H – CI
Thus the reaction involves
Cleavage of one H – H bond, ∆H = 104 kcal
Cleavage of one CI – CI bond, ∆H = 58 kcal
Formation of two H – CI bonds, ∆H = 2 * - 103 kcal
∴ ∆H of the reaction = (104 + 58) – 2(103) = 162 – 206 = - 44 kcal
Now since the enthalpy of formation of a compound is the change in heat content accompanied in the formation of one mole of the compound, the enthalpy of formation of
HCI gas = - 44/2 = - 22 kcal